Question 8 3 pts Given h (32) = (x - 1)2 and g(x) = 3x + 5, find (h+g) (-2). Leave and exact answer.
Answer #1

Given,h(x)= (x-1)²= x²+1-2x

g(x)=3x+5

(h+g)(x)= x²+1-2x+3x+5= x²+x+6.

(h+g)(-2)=(-2)²+(-2)+6=4-2+6=8.

The answer of (h+g)(-2) is 8.